OpenStudy (anonymous):

F(x) = x^-1, x=.9. Use linear approximation to estimate answer

OpenStudy (anonymous):

I get 1+(-1)(1-.9) But that's wrong. PLEASE HELP

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

HELLO MISTY1212!!

OpenStudy (misty1212):

translate as find the equation for the line tangent to the graph of \(y=\frac{1}{x}\) at the point \((1,1)\)

OpenStudy (misty1212):

first take the derivative

OpenStudy (misty1212):

you get \[f'(x)=-\frac{1}{x^2}\]

OpenStudy (misty1212):

then find the slope at \(x=1\) you get \(-1\)

OpenStudy (misty1212):

you with me so far?

OpenStudy (anonymous):

Yep! I think I did all of that correct

OpenStudy (misty1212):

you get \[y-1=-(x-1)\]

OpenStudy (misty1212):

or \[y=-(x-1)+1\]

OpenStudy (misty1212):

now replace \(x\) by \(.9\)

OpenStudy (misty1212):

that is all

OpenStudy (misty1212):

\[-(.9-1)+1=.1+1=1.1\] if my arithmetic is correct

OpenStudy (misty1212):

looks like you switched them somehow

OpenStudy (anonymous):

:-( idk it's saying this is wrong but I think it's completely right. The formula we were given was f(a)+f'(a)(x-a)

OpenStudy (anonymous):

Yeah I switched it but neither is correct on this website hahaha. But it looks right to me

OpenStudy (misty1212):

did you get \(1.1\)?

OpenStudy (misty1212):

i will bet $7 on this answer maybe more

OpenStudy (anonymous):

Pay up!

OpenStudy (anonymous):

Appreciate the help but it's saying it's wrong hahah. I know your work is right. Idk :O

OpenStudy (anonymous):

As defined in Section 3.11, if f(x) is differentiable at x=a, then the approximating function L(x) = f(a) + f'(a)(x-a) is the linearization of f at a. Let f(x)=x^(-1) x0=0.9. Find a linearization at a suitably chosen integer a close to x_0 at which the given function and its derivative are easy to evaluate. L(X) = @misty1212 would reading the question change your answer at all? What did we do wrong here lol

OpenStudy (misty1212):

nothing

OpenStudy (misty1212):

clearly the suitable integer is 1

OpenStudy (misty1212):

also clearly it is \[L(x)=-(x-1)+1\]

OpenStudy (misty1212):

also clearly since \(.9<1\) we should have \(\frac{1}{.9}>1\)

OpenStudy (anonymous):

Ohh it was asking for the formula -(x-1)+1 not for 1.1. Thanks a ton for the help, sorry for the confusion.